Exercise 2-12 Computing Predetermined Overhead Rates and Job Costs [LO2-1, LO2-2, LO2-3] Moody Corporation uses a job-order costing system with a plantwide predetermined overhead rate based on machine-hours. At the beginning of the year, the company made the following estimates Machine-hours required to support estimated production Fixed manufacturing overhead cost Variable manufacturing overhead cost per machine-hour 154,000 656,000 $ 4.70 Required: 1. Compute the plantwide predetermined overhead rate 2. During the year, Job 400 was started and completed. The following information was available with respect to this job Direct materials Direct labor cost Machine-hours used $310 $ 270 39 Compute the total manufacturing cost assigned to Job 400 3. If Job 400 includes 60 units, what is the unit product cost for this job? 4. If Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would it have established for Job 400? Complete this question by entering your answers in the tabs below Required 1 Required 2 Required 3Required 4 Compute the plantwide predetermined overhead rate. (Round your answer to 2 decimal places.) Predetermined overhead rate per MH Required 1 Required 2Required 1 Required 2 Required 3 Required 4 Compute the total manufacturing cost assigned to Job 400. (Round your intermediate calculations to 2 decimal places and your final answer to the nearest dollar amount.) Total manufacturing cost K Required 1 Required 3>Required 1 Required 2 Required 3Required 4 If Job 400 includes 60 units, what is the unit product cost for this job? (Round your intermediate calculations and final answer to the nearest whole dollar.) Unit product cost Required 2 Required 4 )Required 1Required 2 Required 3Required 4 If Moody uses a markup percentage of 120% of its total manufacturing cost, then what selling price per unit would it have established for Job 400? (Round your intermediate calculations and final answer to the nearest whole dollar.) Selling price per unit Required 3 Required 4
Answer #1

1) Predetermine overhead rate = (656000/154000)+4.70 = $8.96 per MH

2) Calculate total manufacturing cost

Direct material 310
Direct labor 270
Manufacturing overhead (39*8.96) 349
Total manufacturing cost 929

3) Unit product cost = 929/60 = $15 per unit

4) Selling price per unit = 15+(15*120%) = $33 per unit